A pharmaccetuical firm maintains that the mean time for a drug to take effect is 24 minutes.In a sample of 400 trails ,the mean time is 26 mintues with a standard deviation of 4minutes.test the hypothesis that the mean time is 24 minute against the alternative that it is not equal to 24 min.use a level of signifixance


1
Expert's answer
2021-01-20T13:23:07-0500

The provided sample mean is "\bar{x}=26" and the sample standard deviation is "s=4."

The sample size is "n=400." The number of degrees of freedom are "df=n-1=399,"

and the significance level is "\alpha=0.01,"


The following null and alternative hypotheses need to be tested:

"H_0: \mu=24"

"H_1: \mu\not =24"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.

Based on the information provided, the critical value for a two-tailed test is "t_c=2.5882."

The rejection region for this two-tailed test is "R=\{t:|t|>2.5882\}"

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{26-24}{4/\sqrt{400}}=10"

Since it is observed that "|t|=10>2.5882=t_c," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu"  is different than 24, at the 0.01 significance level.

Using the P-value approach: The p-value is "p=0," and since "p=0<0.01," it is concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu"  is different than 24, at the 0.01 significance level.